To prevent dehydration, intravenous (IV) solution is given to hospital patients. Commonly used are solutions containing sodium chloride and glucose. A 150.0-mL IV solution contains 6.60 grams of glucose (CHO). What is the molarity of this solution? The molar mass of glucose is 180.16 g/mol.









1
Expert's answer
2022-06-13T13:11:04-0400

Solution:

The molar mass of glucose (C6H12O6) is 180.16 g/mol

Therefore,

Moles of glucose = (6.60 g C6H12O6) × (1 mol C6H12O6 / 180.16 g C6H12O6) = 0.0366 mol C6H12O6


Liters of solution = (150.0 mL) × (1 L / 1000 mL) = 0.150 L


Molarity = Moles of solute / Liters of solution

Therefore,

Molarity of the solution = (0.0366 mol) / (0.150 L) = 0.244 mol/L = 0.244 M

Molarity of the solution = 0.244 M


Answer: The molarity of the solution is 0.244 M